# Iterated sampling importance resampling with adaptive number of proposals

### v1

@@ 8.2. Asymptotic variance
$\mathrm{var}(P_{\lambda},f)=2\langle f|(I-P_{\lambda})^{-1}\rangle_{\pi}-\langle f|f\rangle_{\pi}>\langle f|f\rangle=\mathrm{var}_{\pi}(f).$
@@ 9. Approximate i-SIR in the case of atoms
We will show how approximation of the i-SIR kernel (Definition 5 in Section 4) is derived in the general case. Let us denote by $\hat{P}_{c}$ the transition probability in Lemma 37 in Appendix A, where $c$ is the holding probability (rejection probability when $\pi$ has no atoms), then it also holds that
(13) $\int\hat{P}_{c}(x,\{x\})\mathrm{d}\pi(x)=\int c+(1-c)\pi(\{x\})\pi(\mathrm{d}x).$
@@
(14) $\begin{split}\psi(\lambda)&=\frac{\int_{\mathbb{S}}[P_{\lambda}(x,\{x\})-\pi(\{x\})]\pi(\mathrm{d}x)}{1-\int_{\mathbb{S}}\pi(\{x\})\pi(x)}\\ &=\frac{\int_{\mathbb{S}_{0}}\epsilon(\lambda,x)\pi(\mathrm{d}x)+\int_{\mathbb{S}_{+}}[P_{\lambda}(x,\{x\})-\pi(\{x\})]\pi(\mathrm{d}x)}{1-\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x)}.\end{split}$
Thus we obtain the following:
@@
    $\mathrm{var}_{\pi}(f)\leq\mathrm{var}(\tilde{Q}_{\lambda},f)\leq\frac{4\hat{w}+\lambda-1}{\lambda-1}\mathrm{var}_{\pi}(f)$
@@
Case (i) follows from Lemma 37 in Appendix A, case (ii) from Lemma 32 and case (iii) from Lemma 25. ∎
We note that Definition 5 and Proposition 6 in Section 4 correspond with Definition 34 and Proposition 35, when Assumption (A2) holds.
@@
Both functions $G_{f}$ and $H_{f}$ are well defined (Lemma 24 and 25) and have similar properties compared to the asymptotic variance $V_{f}$ when (A1) holds, namely they are strictly convex decreasing functions and they all behave between the same bounds (Theorem 4, Proposition 6 and 35). These imply that to some extent we can use $G_{f}$ and $H_{f}$ as approximates of $V_{f}$. Nevertheless the smaller $\hat{w}$ gets, the closer the $G_{f}$ and $H_{f}$ are to the asymptotic variance and when $\hat{w}=1$, they are the same:
@@
$\beta(\lambda)\frac{1}{\lfloor\lambda\rfloor}+(1-\beta(\lambda))\frac{1}{\lfloor\lambda\rfloor+1}=b_{\lambda},\qquad\beta(\lambda)\frac{\lfloor\lambda\rfloor-1}{\lfloor\lambda\rfloor}+(1-\beta(\lambda))\frac{\lfloor\lambda\rfloor}{\lfloor\lambda\rfloor+1}=1-b_{\lambda}.$
Combining these with the definitions and Lemma 37 in Appendix A we get the claim. ∎
Namely if $\pi=q$ almost surely, then Algorithm 2 gives the Markov chain in Lemma 37 in Appendix A.
@@ 10.1. Quality of asymptotic variance approximations in a finite state space
In a finite state space (A1) always holds, and the asymptotic variance of a Markov chain can be calculated in principle using the spectral decomposition of its transition probability (Lemma 39 in Appendix A). In the case of i-SIR, the transition and rejection probabilities are not directly available, and direct calculation is too costly for large state space (and/or large number of proposals). Therefore, we consider two cases:
@@
We consider again the same state space $\mathbb{S}=\{1,2,\ldots,6\}$. The probability mass functions $\pi,q:\mathbb{S}\to[0,1]$ shown in Figure 3(a) are the same as in Experiments 10.1 and 10.1 for states $1$, 22, 33 and $4$, and $\hat{w}$ is same as in Experiment 10.1, but $\pi(6)$ and $q(6)$ are $1000$ times smaller than in Experiment 10.1. We choose $M$ and $m$ as before, hence $Y_{u}^{M}=\{6\}$ and $Y_{l}^{m}=\{3\}$. The results in Figure 3 and Table 3 are similar to Experiment 10.1 for most test functions, but the results for $\hat{k}$ resemble those in Experiment 10.1: as $a$ increases, the suboptimality factor increases.
@@ Acknowledgements
Appendix A Asymptotic variance results
@@
Let us have IID random variables $U_{1},\dots,U_{n}\sim Unif(0,1)$ and $X_{0},Z_{1},\dots,Z_{n}\sim\pi$ that are independent of each other. Let
$X_{n}=X_{n-1}I(U_{n}\leq\epsilon)+Z_{n-1}I(U_{n}>\epsilon)$
@@
Thus $g(X_{n})$ are identically distributed and it holds for all $m\geq 0$ that $\mathbb{E}[g(X_{m})]=0$ and $\mathbb{E}[g(X_{m})^{2}]=1$. By induction we get for $0\leq m<n$, that
@@
$=1+\frac{2}{n}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\epsilon^{j-i}=1+\frac{2}{n}\sum_{i=1}^{n}(n-1)\epsilon^{i}.$
We have by Kronecker
$\sum_{i=1}^{n}\frac{\epsilon^{i}}{n}\xrightarrow{n\to\infty}0.$
@@
Let $P=[P_{i,j}]_{i,j=1,\dots,n}$ be a transition probability matrix, such that $P$ is irreducible, aperiodic and $\pi$-reversible with respect to some probability vector $\pi\in\mathbb{R}^{n}$. Let $(X_{k})_{k=0}^{\infty}$ be a Markov chain in state space $\mathbb{S}=(s_{1},\dots,s_{n})$ with transition probability $P$ and initial distribution $\pi$, then for $f:\mathbb{S}\to\mathbb{R}$ it holds
@@
We obtain the given eigenvalues and eigenvectors by [levin-peres-wilmer, Lemma 12.1, Lemma 12-2]. By these properties $P$ is diagonalizable and
@@
$\sum_{k=0}^{\infty}R^{k}=\mathrm{diag}\Big(\frac{1}{1-\lambda_{1}},\dots,\frac{1}{1-\lambda_{m-1}},1,\frac{1}{1-\lambda_{m+1}},\dots,\frac{1}{1-\lambda_{n}}\Big),$
@@
Let $\mathbb{S}=\{s_{1},\dots,s_{n}\}$, $f:\mathbb{S}\to\mathbb{R}$ and $\lambda>1$. It holds for Algorithm 2 with $\mathrm{supp}(\pi)=\mathbb{S}$, that
@@
In the finite discrete case Assumption (A1) holds, thus Algorithm 2 forms $\pi$-reversible Markov chain by Lemma 26 and $P_{\lambda}(x,\{y\})>0$ for all $x,y\in\mathbb{S}$ by Lemma 48 in Appendix E, hence $P_{\lambda}$ is irreducible and aperiodic and thus the claim follows from Lemma 38. ∎
Appendix B Details about the setup
For this weight function holds $w\in[0,\infty)$. We say that $x\in\mathbb{X}$ is an atom for measure $\mu$, when $\mu(\{x\})>0$. For $\mathbb{Q}$ this condition is equivalent to $q(x)>0$. Note that $\mathbb{Q}\ll\mu$ implies that if $x$ is not an atom for $\mu$, then it is neither for $\mathbb{Q}$. However an atom for $\mu$ doesn’t require to be an atom for $\mathbb{Q}$. Similar conditions also hold for $\Pi$ with respect to $\mathbb{Q}$. Note that in general $q(x)=\mathbb{Q}(\{x\})$ doesn’t hold for an atom of $\mathbb{Q}$, but for instance when $\mu$ restricted to the atoms of $\mathbb{Q}$ is counting measure, then $q(x)=\mathbb{Q}(\{x\})$. Especially when $\mu$ is Lebesgue measure/counting measure on $\mathbb{R}$ we can have the usual probability density functions/probability mass functions as the Radon-Nikodym derivatives of $\pi$ and $q$.
@@ Appendix C Results about real sequences and their sums and differences
where equality holds if and only if $(a_{1},\dots,a_{n})$ and $(b_{1},\dots,b_{n})$ are linearly dependent.
@@
$\frac{1}{n}\sum_{k\in J}\frac{\big(\sum_{i=1,i\neq k}^{n+1}b_{i}a_{i}\big)^{2}}{\sum_{i=1,i\neq k}^{n+1}b_{i}}\geq\frac{\big(\sum_{i=1}^{n+1}b_{i}a_{i}\big)^{2}}{\sum_{j=1}^{n+1}b_{j}},$
@@
Assume $A$ and $B$ are linearly dependent i.e. $cA+dB=0$ for $c\neq 0$ and $d\neq 0$, then by (17)
$0=\sum_{k=1}^{n+1}(cA_{k}+dB_{k})=n\sum_{l=1}^{n+1}b_{l}[ca_{l}+d]$
@@
$b_{k}[ca_{k}+d]=\sum_{l=1}^{n+1}b_{l}[ca_{l}+d]-\sum_{m=1,m\neq k}^{n+1}b_{m}[ca_{m}+d]=-(cA_{k}+dB_{k})=0,$
hence for each $k$ has to hold either $b_{k}=0$ or $ca_{k}+d=0$. We get a contradiction, since $a_{i}\neq a_{j}$ and $b_{i},b_{j}>0$. ∎
@@
$=B_{N+1}A_{N}[b_{N+1}-b_{N+2}]+2b_{N+1}b_{N+2}A_{N}$
$\quad-2B_{N}B_{N+2}a_{N+1}+B_{N}B_{N+1}[a_{N+1}+a_{N+2}]$
$=B_{N+1}A_{N}[b_{N+1}-b_{N+2}]+2[B_{N+1}-B_{N}]b_{N+2}A_{N}$
$\quad-2B_{N}[B_{N+1}+b_{N+2}]a_{N+1}+B_{N}B_{N+1}a_{N+1}+B_{N}B_{N+1}a_{N+2}$
$=B_{N+1}A_{N}[b_{N+1}-b_{N+2}]+2B_{N+1}b_{N+2}A_{N}-2B_{N}b_{N+2}A_{N+1}$
$\quad-B_{N}B_{N+1}[a_{N+1}-a_{N+2}].\qed$
@@ Appendix D Details about modification
In the following lemma the assumption in case (i) is generalisation of sequential convexity and matches this, if $a_{i}=i$ for all $i$.
@@
The function $f$ is continuous and for all $i\in\mathbb{Z}$ holds $f(a_{i})=b_{i}$. Additionally if for all $n\in\mathbb{Z}$ holds
@@
We have clearly that $f$ is continuous on the intervals $a_{i},a_{i+1})$, $f(a_{i})=b_{i}$ and $\lim_{x\uparrow a_{i+1}}f(x)=b_{i+1}$ for all $i\in\mathbb{Z}$, hence $f:\mathbb{R}\to\mathbb{R}$ is continuous.
@@
Thus by induction for any $m\in\mathbb{N}$, $m\geq 2$ it holds that
@@
$\langle w^{i}|v\rangle\leq a_{i}(b_{k+1}-b_{k})+\frac{(a_{i}-a_{k})b_{k+1}-(a_{i}-a_{k+1})b_{k}}{a_{k+1}-a_{k}}(a_{k}-a_{k+1})=(w^{k}|v),$
similarly by (19) we have for all $i<k$ that $\langle w^{i}|v\rangle\leq\langle w^{k}|v\rangle$, which together imply $w^{i}\in A$ for every $i\in\mathbb{Z}$. If $x<f(c_{t})$, then $(c_{t},x)\notin A$, since
@@
We have by convexity that $(y,f(y)),(z,f(z))\in A$ and $((1-t)y+tz,(1-t)f(y)+tf(z))\in A$. Therefore $(1-t)f(y)+tf(z)\geq f((1-t)y+tz)$, which proves convexity. We can obtain concavity similarly.
@@
and $f(x)\geq f(a_{k+1})$ is obvious, thus the claim follows. Similarly we obtain strict inequality if $b_{i}>b_{i+1}$, which proves (ii). We obtain (iii) similarly. Both (iv) and (v) are clear by the definition of $f$. ∎
@@
then for $h:[1,\infty)\to\mathbb{R}$
@@
it holds for all $\lambda\geq 1$, that
@@
$h(\lambda)$ $=(N+1-\lambda)h(N)+(\lambda-N)h(N+1)$

### v2

@@ 8.2. Asymptotic variance
$\mathrm{var}(P_{\lambda},f)=2\langle f|(I-P_{\lambda})^{-1}f\rangle_{\pi}-\langle f|f\rangle_{\pi}>\langle f|f\rangle_{\pi}=\mathrm{var}_{\pi}(f).$
@@ 9. Approximate i-SIR in the case of atoms
We will show how approximation of the i-SIR kernel (Definition 5 in Section 4) is derived in the general case. Let us denote by $\hat{P}_{c}$ the transition probability in Lemma 37 in Appendix B, where $c$ is the holding probability (rejection probability when $\pi$ has no atoms), then it also holds that
(13) $\int\hat{P}_{c}(x,\{x\})\pi(\mathrm{d}x)=\int c+(1-c)\pi(\{x\})\pi(\mathrm{d}x).$
@@
(14) $\begin{split}\psi(\lambda)&=\frac{\int_{\mathbb{S}}[P_{\lambda}(x,\{x\})-\pi(\{x\})]\pi(\mathrm{d}x)}{1-\int_{\mathbb{S}}\pi(\{x\})\pi(\mathrm{d}x)}\\ &=\frac{\int_{\mathbb{S}_{0}}\epsilon(\lambda,x)\pi(\mathrm{d}x)+\int_{\mathbb{S}_{+}}[P_{\lambda}(x,\{x\})-\pi(\{x\})]\pi(\mathrm{d}x)}{1-\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x)}.\end{split}$
Thus, we obtain the following:
@@
    $\mathrm{var}_{\pi}(f)\leq\mathrm{var}(\tilde{Q}_{\lambda},f)\leq\frac{4\hat{w}+\lambda-1}{\lambda-1}\mathrm{var}_{\pi}(f)$.
@@
Case (i) follows from Lemma 37 in Appendix B, case (ii) from Lemma 32 and case (iii) from Lemma 25. ∎
We note that Definition 5 and Proposition 6 in Section 4 correspond with Definition 34 and Proposition 35 when Assumption (A2) holds.
@@
Both functions $G_{f}$ and $H_{f}$ are well defined (Lemma 24 and 25) and have similar properties compared to the asymptotic variance $V_{f}$ when (A1) holds, namely they are strictly convex decreasing functions and they all behave between the same bounds (Theorem 4, Proposition 6 and 35). These imply that to some extent we can use $G_{f}$ and $H_{f}$ as approximates of $V_{f}$. Nevertheless, the smaller $\hat{w}$ gets, the closer $G_{f}$ and $H_{f}$ are to the asymptotic variance, and when $\hat{w}=1$, they are the same:
@@
$\beta(\lambda)\frac{1}{\lfloor\lambda\rfloor}+\big(1-\beta(\lambda)\big)\frac{1}{\lfloor\lambda\rfloor+1}=b_{\lambda},\qquad\beta(\lambda)\frac{\lfloor\lambda\rfloor-1}{\lfloor\lambda\rfloor}+\big(1-\beta(\lambda)\big)\frac{\lfloor\lambda\rfloor}{\lfloor\lambda\rfloor+1}=1-b_{\lambda}.$
Combining these with the definitions and Lemma 37 in Appendix B, we obtain the claim. ∎
Namely if $\pi=q$ almost surely, then Algorithm 2 gives the Markov chain in Lemma 37 in Appendix B.
@@ 10.1. Quality of asymptotic variance approximations in a finite state space
In a finite state space (A1) always holds, and the asymptotic variance of a Markov chain can be calculated in principle using the spectral decomposition of its transition probability (Lemma 39 in Appendix B). In the case of i-SIR, the transition and rejection probabilities are not directly available, and direct calculation is too costly for large state space (and/or large number of proposals). Therefore, we consider two cases:
@@
We consider again the same state space $\mathbb{S}=\{1,2,\ldots,6\}$. The probability mass functions $\pi,q:\mathbb{S}\to[0,1]$ shown in Figure 3(a) are the same as in Experiments 10.1 and 10.1 for states $1$ to $4$, and $\hat{w}$ is same as in Experiment 10.1, but $\pi(6)$ and $q(6)$ are $1000$ times smaller than in Experiment 10.1. We choose $M$ and $m$ as before, hence $Y_{u}^{M}=\{6\}$ and $Y_{l}^{m}=\{3\}$. The results in Figure 3 and Table 3 are similar to Experiment 10.1 for most test functions, but the results for $\hat{k}$ resemble those in Experiment 10.1: as $a$ increases, the suboptimality factor increases.
@@ Acknowledgements
Appendix A Details about the setup
For this weight function holds $w\in[0,\infty)$. We say that $x\in\mathbb{X}$ is an atom for measure $\mu$ when $\mu(\{x\})>0$. For $\mathbb{Q}$ this condition is equivalent to $q(x)>0$. Note that $\mathbb{Q}\ll\mu$ implies that if $x$ is not an atom for $\mu$, then it is neither for $\mathbb{Q}$. However an atom for $\mu$ doesn’t require to be an atom for $\mathbb{Q}$. Similar conditions also hold for $\Pi$ with respect to $\mathbb{Q}$. Note that in general $q(x)=\mathbb{Q}(\{x\})$ doesn’t hold for an atom of $\mathbb{Q}$, but for instance when $\mu$ restricted to the atoms of $\mathbb{Q}$ is counting measure, then $q(x)=\mathbb{Q}(\{x\})$. Especially when $\mu$ is Lebesgue measure/counting measure on $\mathbb{R}$ we can have the usual probability density functions/probability mass functions as the Radon-Nikodym derivatives of $\pi$ and $q$.
Appendix B Asymptotic variance results
@@
Let us have i.i.d. random variables $U_{1},\dots,U_{n}\sim Unif(0,1)$ and $X_{0},Z_{1},\dots,Z_{n}\sim\pi$ that are independent of each other. Let
$X_{n}=X_{n-1}I(U_{n}\leq\epsilon)+Z_{n}I(U_{n}>\epsilon)$
@@
hence $g(X_{n})$ are identically distributed and it holds for all $m\geq 0$ that $\mathbb{E}[g(X_{m})]=0$ and $\mathbb{E}[g(X_{m})^{2}]=1$. By induction we get for $0\leq m<n$, that
@@
$=1+\frac{2}{n}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\epsilon^{j-i}=1+\frac{2}{n}\sum_{i=1}^{n}(n-i)\epsilon^{i}.$
We have $\sum_{i=1}^{\infty}\epsilon^{i}=\frac{\epsilon}{1-\epsilon}$, hence by Kronecker
$\sum_{i=1}^{n}\frac{i\epsilon^{i}}{n}\xrightarrow{n\to\infty}0.$
@@
Let $P=[P_{i,j}]_{i,j=1,\dots,n}$ be a transition probability matrix, such that $P$ is irreducible, aperiodic and $\pi$-reversible with respect to some probability vector $\pi\in\mathbb{R}^{n}$. Let $(X_{k})_{k=0}^{\infty}$ be a Markov chain in state space $\mathbb{S}=(s_{1},\dots,s_{n})$ with transition probability $P$ and initial distribution $\pi$, then it holds for $f:\mathbb{S}\to\mathbb{R}$ that
@@
We obtain the given eigenvalues and eigenvectors by [levin-peres-wilmer, Lemma 12.1, Lemma 12.2]. By these properties $P$ is diagonalizable and
@@
$\sum_{k=0}^{\infty}R^{k}=\mathrm{diag}\Big(\frac{1}{1-\lambda_{1}},\dots,\frac{1}{1-\lambda_{m-1}},0,\frac{1}{1-\lambda_{m+1}},\dots,\frac{1}{1-\lambda_{n}}\Big),$
@@
Let $\mathbb{S}=\{s_{1},\dots,s_{n}\}$, $f:\mathbb{S}\to\mathbb{R}$ and $\lambda>1$. It holds for Algorithm 2 with $\mathrm{supp}(\pi)=\mathbb{S}$ that
@@
In the finite discrete case Assumption (A1) holds. Algorithm 2 forms $\pi$-reversible Markov chain by Lemma 26, and $P_{\lambda}(x,\{y\})>0$ for all $x,y\in\mathbb{S}$ by Lemma 49 in Appendix E, hence $P_{\lambda}$ is irreducible and aperiodic. Therefore, the claim follows from Lemma 38. ∎
@@ Appendix C Results about real sequences and their sums and differences
where the equality holds if and only if $(a_{1},\dots,a_{n})$ and $(b_{1},\dots,b_{n})$ are linearly dependent.
@@
$\frac{1}{n}\sum_{k\in J}\frac{\big(\sum_{i=1,i\neq k}^{n+1}b_{i}a_{i}\big)^{2}}{\sum_{j=1,j\neq k}^{n+1}b_{j}}\geq\frac{\big(\sum_{i=1}^{n+1}b_{i}a_{i}\big)^{2}}{\sum_{j=1}^{n+1}b_{j}},$
@@
Assume $A$ and $B$ are linearly dependent i.e. $cA+B=0$ for some $c\neq 0$, then by (17)
$0=\sum_{k=1}^{n+1}(cA_{k}+B_{k})=n\sum_{l=1}^{n+1}b_{l}(ca_{l}+1)$
@@
$b_{k}(ca_{k}+1)=\sum_{l=1}^{n+1}b_{l}(ca_{l}+1)-\sum_{m=1,m\neq k}^{n+1}b_{m}(ca_{m}+1)=-(cA_{k}+B_{k})=0,$
hence for each $k$ has to hold either $b_{k}=0$ or $ca_{k}+1=0$. Since $a_{i}\neq a_{j}$ and $b_{i},b_{j}>0$, we get a contradiction. ∎
@@
$=B_{N+1}A_{N}(b_{N+1}-b_{N+2})+2b_{N+1}b_{N+2}A_{N}$
$\quad-2B_{N}B_{N+2}a_{N+1}+B_{N}B_{N+1}(a_{N+1}+a_{N+2})$
$=B_{N+1}A_{N}(b_{N+1}-b_{N+2})+2(B_{N+1}-B_{N})b_{N+2}A_{N}$
$\quad-2B_{N}(B_{N+1}+b_{N+2})a_{N+1}+B_{N}B_{N+1}a_{N+1}+B_{N}B_{N+1}a_{N+2}$
$=B_{N+1}A_{N}(b_{N+1}-b_{N+2})+2B_{N+1}b_{N+2}A_{N}-2B_{N}b_{N+2}A_{N+1}$
$\quad-B_{N}B_{N+1}(a_{N+1}-a_{N+2}).\qed$
@@ Appendix D Details about modification
In the following lemma the assumption in case (i) is generalisation of sequential convexity and matches this if $a_{i}=i$ for all $i$.
@@
The function $f$ is continuous and for all $i\in\mathbb{Z}$ holds $f(a_{i})=b_{i}$. Additionally, if for all $n\in\mathbb{Z}$ holds
@@
Clearly $f$ is continuous on the intervals $a_{i},a_{i+1})$, $f(a_{i})=b_{i}$ and $\lim_{x\uparrow a_{i+1}}f(x)=b_{i+1}$ for all $i\in\mathbb{Z}$, hence $f:\mathbb{R}\to\mathbb{R}$ is continuous.
@@
Thus, by induction for any $m\in\mathbb{N}$, $m\geq 2$ it holds that
@@
$\langle w^{i}|v\rangle\leq a_{i}(b_{k+1}-b_{k})+\frac{(a_{i}-a_{k})b_{k+1}-(a_{i}-a_{k+1})b_{k}}{a_{k+1}-a_{k}}(a_{k}-a_{k+1})=(w^{k}|v).$
Similarly, by (19) we have for all $i<k$ that $\langle w^{i}|v\rangle\leq\langle w^{k}|v\rangle$. Together these imply $w^{i}\in A$ for every $i\in\mathbb{Z}$. If $x<f(c_{t})$, then $(c_{t},x)\notin A$, since
@@
Since $A$ is convex, we have $(y,f(y)),(z,f(z))\in A$ and $((1-t)y+tz,(1-t)f(y)+tf(z))\in A$, hence $(1-t)f(y)+tf(z)\geq f((1-t)y+tz)$. Thus, we obtain convexity. Concavity follows similarly.
@@
and $f(x)\geq f(a_{k+1})$ is obvious, thus the claim follows. Similarly, we obtain strict inequality if $b_{i}>b_{i+1}$, which proves (ii). We obtain (iii) similarly. Both (iv) and (v) are clear by the definition of $f$. ∎
@@
then for $h:[1,\infty)\to\mathbb{R}$,
@@
it holds for all $\lambda\geq 1$ that
@@
$h(\lambda)$ $=(N+1-\lambda)g(N)+(\lambda-N)g(N+1)$